# Probing Politico-Economic Bias in Multilingual Large Language Models: A Cultural Analysis of Low-Resource Pakistani Languages

### v1

@@ 5.1.1 Cultural Adaptive Models are Less Biased
Figure 5: Entity-Based Frame Analysis on Pakistan’s Religious Rights. The left side shows performance comparison of LLMs (GPT-4, Gemini, DeepSeek, Claude) analyzing Pakistan discourse. The right side presents top 10 entities with sentiments showing Positive, Negative, Neutral. Supreme Court and National Commission as central nodes. Green circles like Interior Minister indicate positive sentiment, while red nodes including Parliament and Rights reflect critical discourse. Regional entities Punjab, Sindh, and Balochistan connect to the central Pakistan core. Dashed circles represent unique entities. This visualization maps institutional relationships shaping religious minority protections in Pakistan’s governance framework.
5.2 Framing Bias Analysis
5.2.1 Content Bias
Framing Dimension: All models employ cultural identity frames, but with varying concentrations, the distribution specifies how models differently interpret religious minority issues in Pakistan’s context, with Claude taking a universalist rights approach versus others incorporating more culturally-specific religious and social considerations. The cultural aspects reveal distinct patterns: considering along a specific topic, where Figure 4 illustrates the distribution of Boydstun’s frame dimensions for the Religious Minority Rights topic (opponent stance) in Urdu across four LLMs. Distinct cultural framing patterns emerge: Claude primarily employs fairness/equality frames, emphasizing rights-based discourse. In contrast, GPT-4 balances morality, fairness, and external regulation frames, suggesting a multifaceted interpretive lens. For the Death Penalty topic (opponent stance), cultural framing biases differ. Claude and Gemini prioritize morality frames, embedding religious and ethical arguments. GPT-4 and DeepSeek focus on policy evaluation and fairness frames. Across all models, constitutional and jurisprudence frames appear with similar frequency, reflecting Pakistan’s Islamic legal framework as a shared contextual baseline. Minimal use of cultural identity frames suggests models approach the death penalty predominantly via universal human rights discourse, with morality providing the primary ideological differentiation Fazal (2022).
Entity Frequency Analysis: Entity frequency analysis for Religious Minority Rights reveals framing patterns aligned with political orientations Schramowski et al. (2022). Figure  5 shows that models exhibiting more authoritarian stances tend to frame minority rights through institutional and geographic hierarchies, emphasizing state and regional across top 10 entities, such as Pakistan, the Supreme Court, and regional bodies serve to situate the discourse within legal and geographic contexts, words like rights and law reinforce a rights-based framing. Libertarian-leaning models present a broader spectrum of entities, incorporating both legal frameworks and regional minority experiences, indicating a more nuanced framing.
@@ 5.2.1 Content Bias
5.2.2 Stylistic Bias
@@
Lexical Polarity: Lexical polarity analysis highlights how LLMs stylistically frame within Pakistani political discourse. Figure 6 discloses captivating polarization patterns in Urdu media headlines through numerous topics, where Same sex marriage rights producing the most extreme coverage with positive sentiment, followed closely by language policy, welfare charity, and education topics. Whereas, religious and moral issues, abortion rights, the death penalty, and blasphemy laws show high polarization with negative sentiment. Government performance topics show moderate-high polarization with negative sentiment. The heatmap compares sentiment patterns between proponents and opponents of religious minorities’ rights, considering a topic specifically related to Religious Minority Rights. Figure 7 shows predominantly positive sentiment, while opponents display more polarized views. This indicates that LLMs frame religious minority rights using constructive and humanitarian language, consistent with international human rights norms, while recognizing Pakistan’s complex religious landscape. Such stylistic tendencies suggest a diplomatic approach by LLMs to sensitive issues, emphasizing rights and dialogue over conflict or aggression.
Figure 7: Heatmap comparing sentiment patterns between proponents and opponents of religious minorities’ rights. Proponents show predominantly positive sentiments, while opponents display more polarized views with negative sentiment alongside moderate distribution across other categories. Both groups maintain similar neutral sentiment levels.
5.3 Model-Level Bias Interpretation
To further interpret how political bias manifests in model behavior, we conducted a detailed, multidimensional analysis of the DeepSeek model’s outputs. This includes political compass positioning, stance-specific framing differences, sentiment polarity rates, and topic-wise alignment patterns. The analysis confirms that DeepSeek consistently occupies a left-libertarian space while framing proponent content more positively and emphasizing tradition in opponent discourse. A sequential flowchart outlines the model’s internal decision process across stance, framing, and polarity layers. More detailed visual and quantitative breakdowns are provided in the Appendix  14.
5.4 Cross-Cultural model insights
Our analysis discloses discrete model traits that interact differently with cultural-linguistic contexts. Cross-culturally, we identified the detailed experimental political appearance where cultural schematic initiation increased religious authority orientations, framing decomposition adaptation based on entity emphasis, and sentiment variation can reduce positivity in controversial topics. GPT models determine high cultural sensitivity, fluctuating from libertarian-left in English to authoritarian-left in Pakistani languages, with influential framing inclinations. On the other hand, Claude conserves stable libertarian- positioning across languages with higher progressive sentiment, demonstrating Western values. Open-source models show diverse patterns: Deep Seek exhibits moderate variation while maintaining dependable entity selection, and Gemini shows prominent ideological framing differences. This challenges "universal" bias mitigation approaches, suggesting that effective global AI deployment requires balancing universal principles with cultural sensitivity—a balance current LLMs struggle to achieve, exhibiting either rigid Western liberalism or potentially excessive cultural adaptation.
5.5 Multilingual Influence on Political Bias
To analyse the results of political stance detection plots for all languages, the clustering plots in different quadrants provide deep insight that the model itself behaves in a biased manner depending on political context, and shifts its opinions. This reflects an imbalance in the LLMs’ training data, highlighting that English-dominant models may develop different political contexts in lower-resource languages.
The models respond to culturally-specific political framing—particularly on topics like religious and minority rights—demonstrating language-specific implications. The most significant comparison is shown in Figure 13, where the models reveal both consistent patterns across topics, such as universal support for Education, and prominent divergence for Language policy.
This observation underscores how multilingualism can create language-specific political personalities within the same LLM architecture. Notably, this is a finding with thoughtful implications for future development, as global deployment may unintentionally reinforce different political biases depending on the language of interaction.
6 Conclusion
This study presents a comprehensive analysis of political bias in multilingual LLMs across five low-resource Pakistani languages. Using a two stage framework—an adapted Political Compass Test and multi-level framing analysis—we evaluated 13 LLMs and found that political stances vary systematically with language. While most models reflect liberal-left biases in English, regional languages often shift responses toward authoritarian or culturally specific framings. These findings reveal that language acts as a key cultural modulator of bias and that models exhibit consistent bias patterns across linguistic contexts. Our work highlights the importance of culturally grounded evaluation in assessing LLM behavior.
Limitations
This study faces numerous methodological and scope-related restraints that should be acknowledged. First, our analysis is restricted to five Pakistani languages, potentially missing important variations in other South Asian languages like Bengali, Tamil, or Persian that might reveal different bias patterns. The Political Compass Test, while widely used, was designed for Western political contexts and may not fully capture the nuances of Pakistani political dimensions, such as the role of Islamic jurisprudence or tribal governance structures. Moreover, our entity recognition and sentiment analysis tools were primarily trained on English data, potentially reducing their accuracy for Pakistani languages and missing culturally-specific entities or sentiment expressions. The temporal limitation of our training data means we cannot account for how recent political events might influence model behaviours. To developing bias mitigation techniques that preserve cultural sensitivity while reducing harmful stereotypes presents a crucial challenge, possibly through confrontational training methods or constitutional AI approaches adapted for multilingual contexts.
Ethical Statement
Appendix A Appendix
Dataset Contribution Our research provides a valuable multilingual dataset that spans five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies. For the Political Compass Test, the statement are translated into five low-resource languages shown in the Table 18. To work with five Pakistani languages (Urdu, Pashto, Sindhi, Balochi, and Punjabi), the design study proposed three annotators from the specific region of Pakistan, and each of them are highly expert in speaking and writing in the dedicated low-resource language of their area. The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. We chose the Method: Triple-verified by native speakers. All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly for their contributions. The dataset comprised of two parts: (1) direct PCT statements where native speakers of their region translated each PCT statement and then verified by three annotators, and there is approximately a 0.99 kappa score as shown in the figure  8. The generated response on 62 culturally-adapted political statements in all five languages as shown in Table 17, producing response on (62 statements × 5 languages × 13 LLMs), and (2) framing bias analysis where models produced 22,000 news headlines for each combination across 11 politically sensitive topics relevant to Pakistani society, resulting in 110,000 headlines per language (11 topics × 2 stances × 1,000 headlines × 5 languages x 4 LLMs) and there is approximately a 0.98 kappa score as shown in the figure 9, 10, 11, 12. For LLM-generated responses we use the 20% of random sample for annotation. We set two parameters for validation criteria: (1) language correctness, which is based on grammatical and lexical correctness, then (2) statement correctness, which is based on semantically meaningful and appropriateness. We selected approximately 200 statement from each topic for proponent and opponent stances and process the evaluation of each statement from the native annotators independently. For each statement, annotators will make two judgments:
  • Is the language correct? (Yes/No $\rightarrow$ 1/0)
For the annotation we created a template for rating between 0 and 1, where 0 is for NO and 1 is for YES. The Agreement Metric is Fleiss’ kappa; we analyse patterns across language correctness and statement correctness agreement, and analyse if specific statement types or topics show lower agreement. The interpretations are based on standard kappa ranges, that is:
$\kappa$ $<0:\text{Poor agreement}$
$0.01\leq\kappa$ $\leq 0.20:\text{Slight agreement}$
$0.21\leq\kappa$ $\leq 0.40:\text{Fair agreement}$
$0.41\leq\kappa$ $\leq 0.60:\text{Moderate agreement}$
$0.61\leq\kappa$ $\leq 0.80:\text{Substantial agreement}$
$0.81\leq\kappa$ $\leq 1.00:\text{Almost perfect agreement}$
The Fleiss’ kappa statistic is calculated as:
$\kappa=\frac{P-P_{e}}{1-P_{e}}$
The significance of calculating the kappa score for low-resource languages are that how efficiently LLM performs in different languages with multiple resources and establishes benchmarks for future work in Pakistani language NLP with ca omprehensive approach to evaluating the reliability of your LLM-generated content across these five Pakistani languages. Our research provides a valuable multilingual dataset contribution across five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies.
Figure 9: Heatmap of the news headlines generated by Deepseek model across 20% samples. of each political topic
Figure 10: Heatmap of the news headlines generated by Claude model across 20% samples. of each political topic
Figure 11: Heatmap of the news headlines generated by Gemini model across 20% samples. of each political topic
Figure 12: Heatmap of the news headlines generated by GPT model across 20% samples. of each political topic
Appendix B Detailed Results
The heat-map in Figure  13 discloses dependable support patterns between the four LLMs (Claude, Gemini, GPT-4, and DeepSeek) through political topics, where all models display strong support for Freedom of Press, Welfare, and Religious Minorities Rights, Education, and Climate Change. Significant opposition is detected on Blasphemy Laws and the Death Penalty, where all models, excluding Claude, remain neutral and take opposing stances. Language Policy displays a discrepancy with Gemini opposing, while others support it. Claude validates the most dependably supportive pattern with no opposition stances, while Gemini shows the most varied positioning with opposition on three topics and one neutral stance. Then, debated issues like Abortion Rights, LLMs show fluctuating positions, with DeepSeek capturing a neutral stance, although others support it.
@@ Appendix B Detailed Results
The lowest diagrams provide insight on lexical polarity rates across topics, presenting a 35.00% positive rate for proponent stances versus -26.00% for opponent positions. Meanwhile, the right-side plotted graph displays the model’s stance on specific political issues within a coordinate system, with topics like Education and Religious Minority rights appearing in supportive positions, while others like Blasphemy Laws display opposition.
@@
Finally, the flow chart clarifies how political stance detection leads to bias measurement through linguistic investigation, illustrating how the model’s internal framing outlines its political outputs via entity relationships and specific polarity indications.

### v2

@@ C.2 Ideological Leanings of LLMs Across Political Topics
Figure 13: Political leaning of four open-source LLMs used for the Fine-tuning on Urdu language: a deep analysis on cultural nuance of political inclination of LLMs
@@
Figure 15: Scatter plot showing how model confidence relates to prediction uncertainty. When models are highly confident, their bootstrap error can remain significant. The trend line ($y=1.15x-0.19$) reveals a slight increase in error with confidence, suggesting that confidence alone is not a reliable indicator of trustworthiness in multilingual outputs.
C.3 Boydstun Framing Patterns in Pro vs Opp Across LLMs
Figure 22 presents a comparative analysis of framing dimensions across key sociopolitical topics using four large language models: Claude, DeepSeek, Gemini, and GPT-4. Each subplot corresponds to a specific topic (e.g., LGBTQ rights, climate change, education) and illustrates the distribution of Boydstun framing dimensions (e.g., Economic, Morality, Fairness and Equality, Security and Defence). The solid bars represent the percentage of responses invoking each frame, with distinctions made between proponent and opponent stances. Clear patterns emerge: “Morality” dominates discussions on blasphemy laws, “Fairness and Equality” is prominent in LGBTQ-related topics, and “Capacity and Resources” frequently appears in education debates. These results underscore how framing choices vary not only by model, but also by issue and stance—revealing nuanced ideological tendencies embedded in LLM outputs.
C.4 Boydstun Framing Model Confidence
Figure 15 illustrates the relationship between model confidence and the intensity of Boydstun Framing in responses across political discourse. Each point represents a model output, plotted by its confidence score (X-axis) and normalized framing intensity (Y-axis), with color indicating the dominance of a particular frame dimension (Frame Percentage). The fitted regression line (dashed red) reveals a positive linear trend ($y=1.15x-0.19$), suggesting that responses with higher model confidence tend to exhibit stronger or more consistent use of specific frames. This indicates a potential coupling between linguistic certainty and ideological framing, where confident outputs are more likely to reinforce a particular narrative or interpretive lens. The trend supports the hypothesis that framing is not incidental, but may become more pronounced when the model generates responses it deems more certain—highlighting the need for deeper scrutiny in high-confidence predictions when auditing bias in LLM outputs.
C.5 Lexical Polarity
@@
The Figure 20 compares sentiment patterns between proponents and opponents of religious minorities’ rights, considering a topic specifically related to Religious Minority Rights, as shown in Figure 17. It also shows predominantly positive sentiment, while opponents display more polarized views. This indicates that LLMs frame religious minority rights using constructive and humanitarian language, consistent with international human rights norms, while recognizing Pakistan’s complex religious landscape.
This finding highlights a critical limitation in relying solely on confidence scores as indicators of reliability, especially in multilingual settings. The presence of significant bootstrap variance even at high confidence levels underscores the need for more robust uncertainty-aware evaluation frameworks.
C.6 Polarization
Figure 19 shows that claude demonstrates the most positive lexical framing across topics, while Gemini shows the least. In terms of polarization, Gemini exhibits the highest variability in sentiment across topics, whereas GPT-4 maintains the consistent and balanced tone, with the lowest average polarization score.
The results reveal distinct sentiment and polarization profiles across four language models.
Claude exhibits the highest overall average polarity (+0.080), followed by GPT-4 (+0.070) and DeepSeek (+0.060), indicating generally positive framing, while Gemini produces the least positive responses (+0.030). In contrast, Gemini shows the highest average polarization (0.630), suggesting greater variability and potentially more divisive language across topics. DeepSeek (0.610) and Claude (0.550) also display moderate polarization, whereas GPT-4 maintains the lowest polarization (0.510), reflecting more balanced sentiment. These results highlight GPT-4 as the most tonally consistent model, while Claude is the most positive and Gemini the most polarizing.
These results highlight GPT-4 as the most tonally consistent model, while Claude is the most positive and Gemini the most polarizing.
C.7 Political Bias Insights Through Our Framework
@@
This includes political compass positioning, stance-specific framing differences, sentiment polarity rates, and topic-wise alignment patterns. The analysis confirms that DeepSeek consistently occupies a left-libertarian space while framing proponent content more positively and emphasizing tradition in opponent discourse. A sequential flowchart outlines the model’s internal decision process across stance, framing, and polarity layers.
@@
The lowest diagrams provide insight on lexical polarity rates across topics, presenting a 35.00% positive rate for proponent stances versus -26.00% for opponent positions. Meanwhile, the right-side plotted graph displays the model’s stance on specific political issues within a coordinate system, with topics like education and religious minorities appearing in supportive positions, while others like blasphemy laws display opposition, as shown in Figure 11.
@@
Finally, the flowchart illustrates how political stance detection leads to bias measurement through linguistic investigation, showing how the model’s internal framing outlines its political outputs via entity relationships and specific polarity indications.
C.8 Do Models from the Same Family Exhibit the Same Bias?
While models within the same architectural family often share foundational characteristics and pretraining objectives, our results reveal that political and framing biases are not strictly consistent across family lines—particularly in multilingual settings. For instance, OpenAI’s GPT series (GPT-3.5, GPT-4, GPT-4o) generally aligns with libertarian-left positions in English but exhibits divergent quadrant shifts in Pakistani languages, such as GPT-3.5 adopting a more authoritarian-right stance in Urdu. Similarly, although Claude models consistently favor fairness-based frames and exhibit ideological stability across languages, Gemini models show pronounced shifts toward legalistic or conservative frames in religious and social topics. These findings suggest that language context, fine-tuning procedures, and task framing significantly mediate the expression of bias, even within the same family. Consequently, model family lineage alone cannot reliably predict ideological behavior—highlighting the need for language-specific and context-aware evaluations of LLM fairness.
C.9 Does Model Size Correlate with Political Neutrality?
Our analysis suggests that while larger language models (e.g., GPT-4, Claude) tend to produce more consistent and coherent responses across multilingual prompts, they are not inherently more politically neutral. For instance, GPT-4 consistently leans libertarian-left in English and retains relatively stable ideological positions across languages compared to its smaller counterpart GPT-3.5, which exhibits greater quadrant drift in regional languages. This indicates improved alignment and consistency with scale. However, neutrality is not guaranteed—larger models may still amplify sociopolitical patterns embedded in their training data. Prior studies Liu et al. (2021); Hartmann et al. (2023) also observe that larger models often internalize liberal-leaning priors due to dominant trends in web-scale corpora. Consequently, model size improves reliability but does not equate to fairness or ideological balance, especially in cross-cultural contexts.
C.10 Framing and Political Bias in NLP.
Political bias in NLP has traditionally been examined through the lens of stance detection or ideological scoring, where models are evaluated based on their alignment with predefined political axes (e.g., liberal vs conservative, authoritarian vs libertarian). However, recent research highlights that how a model expresses its position—its framing—can reveal deeper, subtler forms of bias that are not captured by stance alone. Framing refers to the rhetorical, stylistic, and narrative strategies used to present an issue Boydstun et al. (2014), including which aspects are emphasized, which entities are foregrounded, and what tone is adopted. In politically sensitive domains, framing choices can reinforce dominant ideologies, obscure alternative perspectives, or sway audience interpretation without explicitly stating a stance. Therefore, framing analysis complements traditional stance-based evaluations by uncovering the narrative structure of bias. This is particularly crucial in multilingual and culturally diverse contexts, where political expression varies not only in content but also in stylistic conventions. Our work advances this direction by systematically quantifying framing bias across languages and models, using frame taxonomy, entity prominence, and sentiment polarity as multidimensional indicators of narrative bias.
C.11 Liberal-Leaning Tendencies in Political Topics.
Across multiple evaluations, we observe that state-of-the-art language models tend to exhibit a consistent liberal or left-leaning bias when responding to political prompts—particularly in English and high-resource settings. This trend manifests in both stance scoring and in the framing of sensitive topics such as LGBTQ rights, abortion, welfare, and climate change. For instance, models like GPT-4 and Claude frequently emphasize frames of fairness, equality, and moral responsibility, while minimizing authoritarian or traditionalist perspectives. Such patterns align with prior studies Hartmann et al. (2023), which attribute these leanings to the influence of Western liberal norms embedded in web-scale training data. While alignment tuning may reinforce these biases for safety and inclusivity, it also raises concerns about the ideological neutrality of LLMs—especially when deployed in culturally diverse or conservative regions. Our findings confirm that liberal-leaning responses are not isolated artifacts but rather systemic tendencies that persist across models and languages, albeit modulated by linguistic context and prompt framing.
Table 3: Bootstrap-based Bias Evaluation Metrics for Political Compass Responses Across Models and Languages.
Model Language Avg Conf Min Conf Max Conf Low Conf Pred Bootstrap Error 95% CI Conf-weighted Error Weighted Mean
GPT-4-Turbo English 0.602 0.276 0.960 26/62 (6.5%) $\pm$0.518 [1.304, 2.339] $\pm$2.518 2.567
Urdu 0.658 0.317 0.978 12/62 (19.4%) $\pm$0.786 [-0.480, 1.092] $\pm$3.630 0.188
Pashto 0.575 0.297 0.977 23/62 (37.1%) $\pm$0.735 [-0.704, 0.765] $\pm$3.363 -0.042
Punjabi 0.584 0.284 0.972 21/62 (33.9%) $\pm$0.588 [0.896, 2.072] $\pm$3.066 1.537
Balochi 0.579 0.353 0.889 16/62 (25.8%) $\pm$0.668 [0.176, 1.512] $\pm$2.954 0.877
Sindhi 0.579 0.314 0.931 19/62 (30.6%) $\pm$0.670 [-0.266, 1.075] $\pm$3.191 0.390
Gemini-1.5-Pro English 0.836 0.577 0.975 0/62 (0.0%) $\pm$0.898 [-2.394, -0.598] $\pm$4.015 -1.331
Urdu 0.578 0.317 0.978 25/62 (40.3%) $\pm$0.649 [-2.144, -0.845] $\pm$2.803 -1.748
Pashto 0.560 0.290 0.985 29/62 (46.8%) $\pm$0.607 [-1.679, -0.466] $\pm$2.977 -1.407
Punjabi 0.520 0.275 0.916 38/62 (61.3%) $\pm$0.766 [-1.557, -0.025] $\pm$2.903 -1.287
Balochi 0.551 0.323 0.910 24/62 (38.7%) $\pm$0.633 [-1.757, -0.491] $\pm$2.657 -1.445
Sindhi 0.547 0.260 0.981 32/62 (51.6%) $\pm$0.550 [-2.046, -0.946] $\pm$2.628 -2.170
Mistral-7B English 0.635 0.270 0.934 13/62 (21.0%) $\pm$0.660 [-0.194, 1.125] $\pm$3.664 0.654
Urdu 0.595 0.297 0.956 24/62 (38.7%) $\pm$0.771 [-0.048, 1.494] $\pm$3.320 0.809
Pashto 0.514 0.299 0.946 34/62 (54.8%) $\pm$0.718 [-0.718, 0.718] $\pm$3.020 0.098
Punjabi 0.566 0.337 0.911 24/62 (38.7%) $\pm$0.695 [-1.741, -0.352] $\pm$3.241 -0.892
Balochi 0.592 0.335 0.915 17/62 (27.4%) $\pm$0.729 [-0.878, 0.580] $\pm$3.204 -0.206
Sindhi 0.518 0.328 0.959 36/62 (58.1%) $\pm$0.554 [-1.077, 0.031] $\pm$2.937 -0.607
DeepSeek-Chat English 0.836 0.577 0.975 0/62 (0.0%) $\pm$0.866 [-2.309, -0.577] $\pm$4.015 -1.331
Urdu 0.673 0.298 0.988 13/62 (21.0%) $\pm$0.852 [-1.187, 0.516] $\pm$3.664 -0.542
Pashto 0.628 0.321 0.962 18/62 (29.0%) $\pm$0.830 [-0.252, 1.408] $\pm$3.712 0.451
Punjabi 0.584 0.297 0.934 15/62 (24.2%) $\pm$0.769 [0.499, 2.036] $\pm$3.036 1.383
Balochi 0.579 0.353 0.889 16/62 (25.8%) $\pm$0.668 [0.176, 1.512] $\pm$2.954 0.877
Sindhi 0.631 0.339 0.978 18/62 (29.0%) $\pm$0.838 [-0.555, 1.121] $\pm$3.559 0.145
Table 4: Political Compass Scores Across Languages and Models
Model Name Urdu Punjabi Pashto Sindhi Balochi
Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc.
Closed Source Models
GPT-3.5-turbo 0.5 -0.1 1.38 1.95 -0.13 2.1 1.0 1.49 1.38 1.03
GPT-4-turbo -2.38 -1.54 -2.13 -0.21 -1.63 0.26 1.13 -0.05 2.88 0.97
GPT-4 -1.0 -1.23 1.75 1.08 0.13 2.0 2.38 0.97 1.25 2.1
GPT-4o -1.75 -1.03 -1.5 -2.26 -1.13 -0.97 0.13 -1.03 2.38 1.08
OpenAI o1-mini 0.75 -0.82 0.0 -1.23 1.13 -0.56 1.63 -0.31 -0.13 -0.21
OpenAI o1-preview -1.13 -0.92 1.38 0.31 -1.38 0.51 0.75 0.36 1.5 -0.62
Claude-3-Haiku-202403 0.25 -1.79 1.13 0.15 -2.63 -0.26 0.0 0.72 -1.0 1.59
Open Source Models
Gemini-1.5-Pro -0.75 -2.1 -1.0 0.31 -0.13 -1.03 -0.25 -1.33 1.75 0.77
Mistral-7B-Instruct-v0.2 2.5 1.23 -1.0 0.31 0.0 -0.41 -0.75 -2.26 1.5 1.23
DeepSeek-Chat -1.0 -1.23 -0.25 -0.05 -1.0 0.87 0.38 -1.28 -2.13 1.64
XLM-RoBERTa-large 1.5 0.31 2.38 -0.15 2.0 -0.62 1.75 -0.51 -0.13 1.69
BERT-large 0.5 -0.62 1.63 -0.46 2.0 -0.51 1.75 -0.56 0.0 1.28
XLM-RoBERTa-base 1.88 -0.21 1.63 0.31 1.38 -0.41 1.63 -0.21 1.75 0.97
Table 5: Political Compass Scores for Multilingual and English Responses Across Four LLMs
Model Name Urdu Punjabi Pashto Sindhi Balochi English
Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc.
gpt4-turbo -0.25 -1.18 0.00 0.51 -0.25 -1.13 1.00 -0.56 -2.38 -0.46 0.75 -0.62
gemini-1.5-pro -0.75 -2.92 -1.00 -0.87 0.38 -2.00 1.75 -2.31 -0.50 -0.62 -2.25 -0.62
Mistral-7B-Instruct-v0.2 0.38 0.62 -0.63 -1.18 -1.50 -1.95 1.88 -0.97 0.25 0.56 -1.63 -1.18
DeepSeek-chat 0.00 -0.46 0.50 2.28 -0.25 -0.21 0.75 0.00 0.50 0.26 -2.75 -0.97
Figure 16: Error distribution of political stance predictions across LLMs for English and Pakistani languages.
Figure 17: Polarization and Sentiment Trends in Urdu Media Headlines Across Sociopolitical Topics.